# The Dual-Route Model of Induction

### v1

@@ D.4 Qualitative Examples
Box D.3: (Llama-3-8b) Original Model vs. Top-32 Token Induction Heads Ablated I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb I’ve reread my comments on his lines, and found that I’d borrowed from his fiery orb a kind of opalescent light
@@
Box D.4: (OLMo-2-7b) Original Model vs. Top-32 Token Induction Heads Ablated I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb I have read his lines again, and I have not found them without pleasure. In many cases, I have caught myself borrowing a kind of opalescent light from his fiery orb.
@@
Box D.5: (Pythia-6.9b) Original Model vs. Top-32 Token Induction Heads Ablated I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb I have reread my comments on his lines, and I have found them to be, in general, true, and I have found myself to be in agreement with them.
@@
Box D.6: (Pythia-6.9b*) Original vs. Top-16* Token Induction Heads Ablated I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb I have reread, not without pleasure, my comments to his lines, and have caught myself in many cases, borrowing a kind of opalescent light from my poet’s fiery orb
@@
Box D.7: (Llama-2-7b) Original Model vs. Top-32 Token Induction Heads Ablated foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [item for sublist in     bar for item in sublist     if i % 2 == 0]
@@
Box D.8: (OLMo-2-7b) Original Model vs. Top-32 Token Induction Heads Ablated foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [x for x in bar if     x % 2 == 0]
@@
Box D.9: (Pythia-6.9b) Original Model vs. Top-32 Token Induction Heads Ablated foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = foo.append(foo) print(foo)
@@
Box D.10: (Llama-2-7b) Original Model vs. Top-32 Token Induction Heads Ablated English vocab: 1. hospital 2. visage 3. unlucky 4. lesbian 5. know 6. wizard 7. unfair 8. sound 9. cortical 10. obsolete English vocab:
@@ 10. The
Box D.11: (Llama-3-8b) Original Model vs. Top-32 Token Induction Heads Ablated English vocab: 1. hospital 2. visage 3. unlucky 4. lesbian 5. know 6. wizard 7. unfair 8. sound 9. cortical 10. obsolete English vocab:
@@
Box D.12: (OLMo-2-7b) Original Model vs. Top-32 Token Induction Heads Ablated English vocab: 1. hospital 2. visage 3. unlucky 4. lesbian 5. know 6. wizard 7. unfair 8. sound 9. cortical 10. obsolete English vocab:
@@
Box D.13: (Pythia-6.9b) Original Model vs. Top-32 Token Induction Heads Ablated English vocab: 1. hospital 2. visage 3. unlucky 4. lesbian 5. know 6. wizard 7. unfair 8. sound 9. cortical 10. obsolete English vocab:
@@
Appendix E Concept Induction is Language-Agnostic
We show results from Section 5 for more languages, with both Llama models. Figure 36 shows results for the same experiment as in Figure 6, expanding evaluation to Llama-3-8b and two more language sets: patching from French-English into German-Russian, and patching from Russian-Spanish into English-Japanese.
@@ Appendix E Concept Induction is Language-Agnostic
Appendix F Function Vector Versus Concept Induction
@@
Figure 37: Results for patching the top-$k$ FV heads from one translation prompt to another. Top-left is the same plot as Figure 7. Patching FV heads from Spanish-Italian to Japanese-Chinese does not flip the output language to Italian, but patchiing from French-English to German-Russian flips the output language to English.

### v2

@@ D.4 Qualitative Examples
Box D.1: (Llama-3-8b) Original Model vs. Top-32 Token Induction Heads Ablated I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb I’ve reread my comments on his lines, and found that I’d borrowed from his fiery orb a kind of opalescent light
@@
Box D.2: (OLMo-2-7b) Original Model vs. Top-32 Token Induction Heads Ablated I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb I have read his lines again, and I have not found them without pleasure. In many cases, I have caught myself borrowing a kind of opalescent light from his fiery orb.
@@
Box D.3: (Pythia-6.9b) Original Model vs. Top-32 Token Induction Heads Ablated I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb I have reread my comments on his lines, and I have found them to be, in general, true, and I have found myself to be in agreement with them.
@@
Box D.4: (Pythia-6.9b*) Original vs. Top-16* Token Induction Heads Ablated I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb I have reread, not without pleasure, my comments to his lines, and in many cases have caught myself borrowing a kind of opalescent light from my poet’s fiery orb I have reread, not without pleasure, my comments to his lines, and have caught myself in many cases, borrowing a kind of opalescent light from my poet’s fiery orb
@@
Box D.5: (Llama-2-7b) Original Model vs. Top-32 Token Induction Heads Ablated foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [] for i in range(len(bar)):     if i % 2 == 0:         asdf.append(bar[i]) foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [item for sublist in     bar for item in sublist     if i % 2 == 0]
@@
Box D.6: (OLMo-2-7b) Original Model vs. Top-32 Token Induction Heads Ablated foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [] for i in range(len(bar)):     if i % 2 == 0:         asdf.append(bar[i]) foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [x for x in bar if     x % 2 == 0]
@@
Box D.7: (Pythia-6.9b) Original Model vs. Top-32 Token Induction Heads Ablated foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = [] for i in range(len(bar)):     if i % 2 == 0:         asdf.append(bar[i]) foo = [] for i in range(len(bar)):     if i % 2 == 0:         foo.append(bar[i]) foo = foo.append(foo) print(foo)
@@
Box D.8: (Llama-2-7b) Original Model vs. Top-32 Token Induction Heads Ablated English vocab: 1. hospital 2. visage 3. unlucky 4. lesbian 5. know 6. wizard 7. unfair 8. sound 9. cortical 10. obsolete English vocab:
@@ 10. The
Box D.9: (Llama-3-8b) Original Model vs. Top-32 Token Induction Heads Ablated English vocab: 1. hospital 2. visage 3. unlucky 4. lesbian 5. know 6. wizard 7. unfair 8. sound 9. cortical 10. obsolete English vocab:
@@
Box D.10: (OLMo-2-7b) Original Model vs. Top-32 Token Induction Heads Ablated English vocab: 1. hospital 2. visage 3. unlucky 4. lesbian 5. know 6. wizard 7. unfair 8. sound 9. cortical 10. obsolete English vocab:
@@
Box D.11: (Pythia-6.9b) Original Model vs. Top-32 Token Induction Heads Ablated English vocab: 1. hospital 2. visage 3. unlucky 4. lesbian 5. know 6. wizard 7. unfair 8. sound 9. cortical 10. obsolete English vocab:
@@
Appendix E Concept and Token Lens
E.1 Approach
Every attention head in an autoregressive transformer can be thought of as “reading” some small amount of information from all previous token positions and subsequently “writing” that information back into the current residual stream. Can we visualize the semantic information that concept induction heads are contributing to the residual stream?
Let $d$ be a model’s hidden dimension and $m<d$ be the dimension of a single head. We rely on a key insight from Elhage et al. (2021): that the value and output projections for a particular head $h$ at layer $l$, $\smash{V_{(l,h)}\in\mathbb{R}^{(m,d)}}$ and $\smash{O_{(l,h)}\in\mathbb{R}^{(d,m)}}$ respectively, are solely responsible for whatever information a head writes into the residual stream. Specifically, they point out that the product of these two matrices $O_{(l,h)}V_{(l,h)}$ is a low-rank $d\times d$ matrix (at most rank $m$) that determines the effect of head $(l,h)$ on the residual stream. In other words, multiplying a hidden state $x_{l}$ by this matrix extracts whatever information within $x_{l}$ that this head typically contributes to the residual stream.
To build a concept lens $\smash{L_{C_{k}}\in\mathbb{R}^{(d,d)}}$ that reads from all of the concept induction head subspaces simultaneously, we combine the weights from the top-$k$ concept induction heads $C_{k}$. We choose $k=80$ based on results from Section 6, and calculate the sum of all concept OV matrices:
$L_{C_{k}}=\sum_{(l,h)\in C_{k}}V_{(l,h)}O_{(l,h)}.$ (5)
If all attention heads in $C_{k}$ are in the same layer, $L_{C_{k}}x_{l}$ is mathematically equivalent to taking the sum of the outputs of those attention heads. However, we also allow for summation of heads across layers, which was empirically effective in prior work (Todd et al., 2024), possibly because transformer representations are interchangeable in intermediate layers (Lad et al., 2025).
Finally, we can project $L_{C_{k}}x_{l}$ to token space by applying the model’s final normalization module and decoder head (nostalgebraist, 2020). This approach works for any set of heads: we can create a token lens $L_{T_{k}}\in\mathbb{R}^{(d,d)}$ out of the top-$k$ token induction heads, or a baseline lens $L\in\mathbb{R}^{(d,d)}$ as the sum of all OV matrices in the model.
E.2 Lens Output Examples
Figures 39 through 42 show concept lens outputs for three different models. Compared to a baseline lens that sums all OV matrices (Figure 44), these plots reveal differing semantics of the same token in three different contexts. We also show examples of token lens in Figure 43, which appears ineffective for inals (the last token of a multi-token word) but clearly reveals token-level information at other positions. We posit that this discrepancy is a consequence of the “token erasure” effect found by Feucht et al. (2024).
Figure 39: Concept lens outputs for Llama-2-7b. We multiply the hidden state for inals at every layer by $L_{C_{k}}$ (Equation 5) before projecting to vocabulary space. Applying this lens reveals the semantics of the token inals, which depends on the context.
Figure 40: Concept lens outputs for Llama-3-8b. We multiply the hidden state for inals at every layer by $L_{C_{k}}$ (Equation 5) before projecting to vocabulary space. Applying this lens reveals the semantics of the token inals, which depends on the context. Unlike other models, early-middle layers are less decodable than middle-late layers. STL is likely a reference to the St. Louis Cardinals.
Figure 41: Concept lens outputs for OLMo-2-7b. We multiply the hidden state for inals at every layer by $L_{C_{k}}$ (Equation 5) before projecting to vocabulary space. Applying this lens reveals the semantics of the token inals, which depends on the context. Unlike Llama models, we must add “STL” to the leftmost prompt to decode semantics related to sports.
Figure 42: Concept lens outputs for OLMo-2-1b. We choose $k=20$, i.e. the top 8% of heads, to be consistent with larger models. Unlike larger models, the signal provided by concept lens is quite noisy. However, for a simple multi-token concept like “New York”, concept lens still reveals more semantic information than a baseline sum of all OV matrices.
Figure 43: Token lens outputs for Llama-2-7b at three token positions. Applying token lens reveals the token that corresponds to a given hidden state. We multiply the hidden state for inals at every layer by $L_{T_{k}}$ before projecting to vocabulary space, with $k=80$. Token lens is not effective for (inals), likely due to the “token erasure” effect for multi-token words described by Feucht et al. (2024).
Figure 44: Baseline “all heads” lens outputs for Llama-2-7b across three prompts. We multiply the hidden state for inals at every layer by $L$, the sum of all OV matrices in the model, before projecting to vocabulary space. Although we can still observe some semantic information when using all heads, concept lens (Figure 39) provides a cleaner signal.
Appendix F Concept Induction is Language-Agnostic
We show results from Section 6 for more languages, with both Llama models. Figure 45 shows results for the same experiment as in Figure 6, expanding evaluation to Llama-3-8b and two more language sets: patching from French-English into German-Russian, and patching from Russian-Spanish into English-Japanese. Figure 46 shows results for a smaller model from a different family, OLMo-2-1b.
@@ Appendix F Concept Induction is Language-Agnostic
[…]
Appendix G Function Vector Versus Concept Induction
@@
Figure 47: Results for patching the top-$k$ FV heads from one translation prompt to another. Top-left is the same plot as Figure 7. Patching FV heads from Spanish-Italian to Japanese-Chinese does not flip the output language to Italian, but patching from French-English to German-Russian flips the output language to English.